OpenStudy (anonymous):

What is the sum of a 52-term arithmetic sequence where the first term is 6 and the last term is 363?

OpenStudy (anonymous):

since th first term and the last term of the arithmetic sequence as well as the number of terms is given to find sum use S=(n/2)*(firstterm + last term)

OpenStudy (anonymous):

here n=52 first term=6 last term=363 thus reqd sum = 52/2 *(6+363)=26*369 =9594

OpenStudy (anonymous):

thanks